• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:03:01 on September 6,2001

    Entry number 73542

    keyword=swing shift summary

    9/5/01 swing shift (16-24)
    Shift workers: J. P. Chen, C. Perdrisat, C. Butuceanu

    Shifts start with beam tuning.
    16:15 beam position, size check. OTR shows beam size ~ 170 um.
    then raster check (run 3296/31296).
    16:20-16:40 NMR/EPR=41.7/40.4%
    16:40-16:55 pol. He3 run at kin5, out/0degree. Short run, since Hall C wanted to
    do charge asymmetry test.
    16:55-17:15, beam off
    17:15-17:35 EMPTY cell run.
    17:35- 18:21 pol. He3 run (Hall C has not started their charge asymmetry test yet!).
    18:21 right arm deadtime went to 100%, endrun failed, killed coda. ROC15 was stuck. Rebooted ROC15. Restarted, ok.
    18:40-18:45 Tried to start N2 run (hall C started the charge asymmetry test), left DAQ not working, ROC14 stuck. Reboot ROC14, restarted CODA, it's ok now.
    18:45-19:21 N2 run (3303/31301).
    19:21-19:50 beam off due to tuning for hall c charge asymmetry test.
    19:50 resume production run 3304/31302.
    19:55 left dipole power tripped off. Stopped run 3304/31302.
    20:00-20:45 Right arm data only run 31303, while left diople recovering.
    20:45 Resume both arm data taking.
    22:05 Since charge asymmetry was large, fluctuating between +-200 ppm.
    increased slope from -17 ppm/step to -25 ppm/step to see if the fluctuation will be reduced. However, the parity DAQ automatically overwrite it back to -17 ppm/step.
    22:46 completed kin5, out/0 degree data taking (2.1M).
    22:46-23:05 NMR/EPR=42.3/42.2%,
    23:05-end of shift data taking at Kin5 OUT/270degree.

    Run 3296/31296, raster check
    RUN 3298/31298: EMPTY cell run, use same cut as pol. He3,
    good e=2.97K/1.60K.
    31299 DAQ problem, ednrunfailed, ROC15 stuck.
    3300-3302 and 31300 junks, fixing DAQ problems.
    3303/31301 N2 run, N2 pressure started at 100 psig ended at 86 psig.
    Use same cut as pol. He3, good e =266K/271K.
    3304/31302 pol. He3 short run due to left dipole power trip.
    31303 pol He3 right arm only.
    31304 just to get run numbers in sync.

    Good e for kin5 out/0degree before this shift is 1.237M
    L-Run# good e R-Run# Good e Grand total
    3297 14K 31297 14K 1.265M
    3299 98K 31299 87K 1.450M
    31303 87K 1.537M
    3305 143K 31305 143K 1.825M
    3306 159K 31306 159K 2.143M

    Kin5, Out/270degree
    L-Run# good e R-run# good e grand total
    3307/31307 to be completed and analyzed by next shift